OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. (1 point) y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

(y-y1)=m(x-x1) here m =-6 (x1,y1)=(-2,-3) now just put these it in the above equation and show me your works with answer

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

can someone please better explain to me

OpenStudy (anonymous):

However A is the correct answer (y-y1)=m(x-x1) ............ this is formula of point slope form here in the question slope is -6 okay ? we would represent slope by m. so m=-6 okay the point is given in the Question (x1,y1)= (-2,-3) just put this point in the above formula and get the anwser

OpenStudy (anonymous):

y-(-3)=-6(x-(-2)) since - multiply with - = + okay y+3=-6(x+2) answer